OpenStudy (anonymous):

The temperature of a roast varies according to Newton's Law of Cooling: dT dt equals negative k times the quantity T minus A, where T is the water temperature, A is the room temperature, and k is a positive constant. If a room temperature roast cools from 68°F to 25°F in 5 hours at freezer temperature of 20°F, how long (to the nearest hour) will it take the roast to cool to 21°F?

OpenStudy (anonymous):

https://gyazo.com/23b71615e032fca0d654bb5437697556

OpenStudy (anonymous):

@jim_thompson5910 Help always appreciated :)

jimthompson5910 (jim_thompson5910):

were you able to find the T(t) function? T = temperature t = time

OpenStudy (anonymous):

No

jimthompson5910 (jim_thompson5910):

\[\Large \frac{dT}{dt} = -k(T-A)\] \[\Large \frac{dT}{T-A} = -k dt\] \[\Large \int\frac{dT}{T-A} = \int-k dt\] I'll let you finish up with the integration. Tell me what you get

OpenStudy (anonymous):

@jim_thompson5910 12?

jimthompson5910 (jim_thompson5910):

how are you getting 12 ?

OpenStudy (anonymous):

Would I integrate both sides?

jimthompson5910 (jim_thompson5910):

yeah what do you get when you do so

OpenStudy (anonymous):

I'm confused, how would I integrate the left side?

jimthompson5910 (jim_thompson5910):

\[\Large \frac{dT}{dt} = -k(T-A)\] \[\Large \frac{dT}{T-A} = -k dt\] \[\Large \int\frac{dT}{T-A} = \int-k dt\] \[\Large \ln(|T-A|) = -k*t + C\] \[\Large |T-A| = e^{-k*t + C}\] \[\Large |T-A| = e^{-k*t}*e^{C}\] \[\Large |T-A| = e^{C}*e^{-k*t}\] what comes next?

OpenStudy (anonymous):

Having two equations, one positive and one negative.

OpenStudy (baru):

T: roast temp (25 deg after 5 hrs) A: freezer temp (20 deg) so T-A is positive, so you can ignore the modulus sign substitute T,A and t=5

OpenStudy (baru):

you need to find k and c, so you need another equation, so substitute t=0 and T=68

OpenStudy (anonymous):

Final answer is 9